We have two situations:
Assignment 10
All Together, Now
a) Pull 5 cards from a standard deck, replacing and shuffling each time, and count the number of face cards.
b) Deal five cards from a standard deck (without replacement). Count the number of face cards.
1. Create probability distributions for both of these situations. Include your tables. [4]
2. Calculate the expected number of face cards for each situation. [2]
3. Explain which probability distribution type you picked for each, and why. [2, Communication]
1
Expert's answer
2020-07-14T17:51:54-0400

a)

"X -" random value, the number of faced cards

"P(X)\; -" probability of getting "X" faced cards with replacing


3.

Here is a binomial distribution. Experiment: shuffle a deck of cards, pull one card. Experiment repeats 5 times. Success is to pull faced card. So probability to pull "k" faced cards after 5 experiments is "C_5^k \cdot p^k \cdot q^{5-k}" , where "p = 12/52,\; q = 1-p=40/52".


1.

"\def\arraystretch{1.6}
 \begin{array}{c|l}
 X & P(X) \\ \hline
 0 & \left(\frac{40}{52}\right)^5 \approx 0.2693 \\ \hdashline
 1 & 5 \cdot\left(\frac{12}{52}\right)^1\cdot\left(\frac{40}{52}\right)^4 \approx 0.404 \\ \hdashline
 2 & 10\cdot\left(\frac{12}{52}\right)^2\cdot\left(\frac{40}{52}\right)^3 \approx 0.2424 \\ \hdashline
 3 & 10\cdot\left(\frac{12}{52}\right)^3\cdot\left(\frac{40}{52}\right)^2 \approx 0.0727 \\ \hdashline
 4 & 5 \cdot\left(\frac{12}{52}\right)^4\cdot\left(\frac{40}{52}\right)^1 \approx 0.0109 \\ \hdashline
 5 & \left(\frac{12}{52}\right)^5 \approx 0.0007 \\
\end{array}"


2.

"E = 0\cdot0.2693 + 1\cdot0.404 + 2\cdot0.2424 + 3\cdot0.0727+4\cdot0.0109+5\cdot0.0007 \approx 1.154"


b)

"X -" random value, the number of faced cards

"P(X)\;-" probability of getting X faced cards without replacing


3.

Here is a hypergeometric distribution. Selections are made from two groups without replacing members of the groups. Success group consists of 12 elements, second group consists of 40 elements. So probability to pull "k" elements from success group after 5 draws is "C_{12}^k \cdot C_{40}^{5-k}/C_{52}^5" .


1.

"\def\arraystretch{1.6}
 \begin{array}{c|l}
 X & P(X) \\ \hline
 0 & C_{12}^0 \cdot C_{40}^{5} / C_{52}^5 \approx 0.2532 \\ \hdashline
 1 & C_{12}^1 \cdot C_{40}^{4} / C_{52}^5 \approx 0.422 \\ \hdashline
 2 & C_{12}^2 \cdot C_{40}^{3} / C_{52}^5 \approx 0.2509 \\ \hdashline
 3 & C_{12}^3 \cdot C_{40}^{2} / C_{52}^5 \approx 0.066 \\ \hdashline
 4 & C_{12}^4 \cdot C_{40}^{1} / C_{52}^5 \approx 0.0076 \\ \hdashline
 5 & C_{12}^5 \cdot C_{40}^{0} / C_{52}^5 \approx 0.0003 \\
\end{array}"


2.

"E = 0\cdot0.2532 + 1\cdot0.422 + 2\cdot0.2509 + 3\cdot0.066+4\cdot0.0076+5\cdot0.0003 \approx 1.1537"